OpenStudy (lifeisadangerousgame):

The product of two consecutive even integers is 120. How do I write an algebraic expression for each even number and write an algebraic equation for this problem?

OpenStudy (anonymous):

let n be the first integer. what is the second integer if they are consecutive even integers (in terms of n)

OpenStudy (lifeisadangerousgame):

well: x(x + 2)

OpenStudy (anonymous):

good and what does that equal to?

OpenStudy (lifeisadangerousgame):

x^2 + 2x

OpenStudy (anonymous):

and according to the problem, what numerical value is that equal to?

OpenStudy (lifeisadangerousgame):

one of the integers that equal to 120?

OpenStudy (anonymous):

well, what i meant was x^2 + 2x = 120 an equation you can solve

OpenStudy (lifeisadangerousgame):

Oh :P yea, so the answer to those two questions is: Algebraic expression: x(x + 2) and x(x + 2) Algebraic equation: x^2 + 2x = 120 ?

OpenStudy (phi):

expression for each even number: n and (n+2) equation: n(n+2)= 120 or n^2+2n-120=0

OpenStudy (lifeisadangerousgame):

THanks guys!